OpenStudy (123ab456c):

Which measurement is the width of this rectangle? Area = 112 cm² A. 6 cm B. 8 cm C. 10 cm D. 12 cm

OpenStudy (alexandervonhumboldt2):

l*w=area

OpenStudy (alexandervonhumboldt2):

thus wight=area/lenght

OpenStudy (123ab456c):



OpenStudy (alexandervonhumboldt2):

area=112 l=14 w=area/l=112/14=?

OpenStudy (123ab456c):

this is the picture that goes with it

OpenStudy (123ab456c):

wait....... is the answer B

OpenStudy (alexandervonhumboldt2):

this is the rectangle, not a square @Arihangdu

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (alexandervonhumboldt2):

you re correct it is B

OpenStudy (123ab456c):

yesss!!!! thanks you all of you that helped me

OpenStudy (alexandervonhumboldt2):

:)

OpenStudy (arihangdu):

I'm sorry yah it is b

OpenStudy (123ab456c):

its ok!! ;)

OpenStudy (arihangdu):

😏